OpenStudy (anonymous):

Papa's Italian Restaurant has collected data about customer sauce orders. It calculated that P(marinara) = 0.74, P(pesto) = 0.62, and P(marinara or pesto) = 0.90. Determine the P(marinara and pesto). 0.12 0.46 0.16 0.22

OpenStudy (anonymous):

i am math illiterate i would like to know how to solve these

OpenStudy (anonymous):

Basically i just want like a walkthrough on how to solve these

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

im actually gonna write this down brb i need copy paper.

ganeshie8 (ganeshie8):

we can use a direct formula : P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

ok im with ya so far

ganeshie8 (ganeshie8):

` P(marinara) = 0.74, P(pesto) = 0.62, and P(marinara or pesto) = 0.90. Determine the P(marinara and pesto).` P(marinara or pesto) = P(marinara) + P(pesto) - P(marinara and pesto)

ganeshie8 (ganeshie8):

plugin the given values : 0.90 = 0.74 + 0.62 - P(marinara and pesto)

ganeshie8 (ganeshie8):

P(marinara and pesto) = 0.74 + 0.62 - 0.90 simplify

OpenStudy (anonymous):

k hold on 1 sec. ill simplify it

OpenStudy (anonymous):

.46?

OpenStudy (anonymous):

0.46 i mean

ganeshie8 (ganeshie8):

Excellent !

OpenStudy (anonymous):

thanks so much